Fermi National Laboratory
 

Accelerator Update for March 2, 2007 - March 5, 2007

 

Friday March 2

The day shift began with Operations monitoring the Recycler stash (83.9E10), with the Antiproton Source (Pbar) stacking (31.2E10), with operators conducting a search and secure of Tevatron (TeV) sectors D, with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations began preparing CDF for an access at 8:24 AM.

At 8:26 AM, the NuMI beam permit tripped off due to bad beam positions. It reset okay.

At 9:43 AM, the MI beam permit tripped off due to the MI-8 collimators.

Meson experts began foil calibration tests in Meson enclosures M01, 2, and 3 around 8:31 AM. They had finished and were accessing the enclosures about two hours later.

CDF came out of access at 9:57 AM.

Operators reported at 10:22 AM that they found a TeV sector D door not latched and was difficult to latch.

At 10:28 AM, MI tripped off due to the MI-8 collimators for the first of four times in about two hours. It reset each time.

At 11:08 AM, Operations hipotted the TeV and found a possible ground fault. Operators and experts investigated.

At 1:09 PM, Operations reported that the TeV ground fault was in the tunnel.

Operations turned off the Linac beam at 1:14 PM so that Linac technicians could repair a RF station (LRF4).

Operations prepared the TeV for an access at 1:20 PM.

At 2:31 PM, operators reported finding an LCW leak in the power supply switch cabinet at the TeV sector C2 service building. Operations contacted experts.

Operations reported at 3:10 PM that they found the TeV ground fault; a copper line (possibly an air-line) in sector B0 was found too close to a high voltage bus.

Operations reported at 3:11 PM that Linac experts completed their repairs on LRF4 and 5.

Operations reported at 4:33 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations began the TeV turn on at 4:44 PM.

Operations put the TeV through a dry squeeze at 5:15 PM.

Operations put the TeV through a wet squeeze at 7:29 PM.

Operations put the TeV into shot setup at 8:16 PM.

Operations established store 5259 at 10:52 PM with an initial luminosity of 178.94E30.

At 11:37 PM, MI tripped off due to the MI-8 collimators. Everything reset okay.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Saturday March 3

The midnight shift began with Operations monitoring store 5259 (136.93E30) the stash (58.2E10), with Pbar stacking (22.6E10), and with NuMI taking beam.

At 1:09 AM, LRF3 tripped off and would not reset remotely. Operators investigated and reset it locally.

At 6:16 AM, store 5259 aborted due to a possible problem with a quadrupole power supply (T:QDD1). There was no quench. After investigations, experts could not say that QDD1 was the cause of the abort, but just to be safe they tightened all the connections and replaced the phase detector.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

At 10:04 AM, A Booster RF station (BRF6) tripped off. It reset.

Operations put the TeV through a dry squeeze at 10:40 AM.

Operations put the TeV into shot setup at 11:12 AM.

At 12:07 PM, NuMI tripped off due to a quadrupole power supply’s (QD108) current drifting out of tolerance. Everything reset okay.

At 1:02 PM, NuMI tripped off on a horn ground fault indication. It reset.

Operations established store 5261 at 1:26 PM with an initial luminosity of 188.59E30.

At 5:08 PM, NuMI tripped off due to the QF112’s power supply current drift. Everything reset.

At 6:41 PM, a Linac RF station (KRF6) tripped off due to a cavity LCW flow fault for the first of two times in five minutes. It reset okay.

Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours.

At 8:02 PM, KRF6 tripped off three times in about three hours. It reset each time.

At 11:30 PM, NuMI tripped off due to a target fan problem. Operations turned off NuMI beam and contacted experts.

Sunday March 4

The midnight shift began with Operations monitoring store 5261 and the stash, with Pbar stacking, and with NuMI off due to a target fan problem.

At 2:28 AM, Operations prepared MI-65 for access so that NuMI experts could check out their target fan.

At 2:28 AM, Operations reported that the Booster had been tripping due to extraction problems. Operators investigated and reported hearing sparking behind the kicker’s (MKS07) Thyratron cabinet. Operations called in experts.

Booster experts reported at 3:45 AM that they had repaired a bad cable for MKS07. Linac resumed sending beam at 3:46 AM.

NuMI experts reported at 4:37 AM that they had fixed the fan problem.

Operations established beam to NuMI at 4:38 AM.

At 6:06 AM, KRF3 tripped off. It reset.

Operations reported at 7:58 AM that the Recycler had stashed four times during the midnight shift.

Operations terminated store 5261 at 8:03 AM.

Operations put the TeV into shot setup at 8:15 AM.

At 8:41 AM, NuMI tripped off due to the QF112’s power supply drifting low. It reset.

At 9:20 AM, a cryo system helium cold compressor (MYCOM #3) had tripped off at the TeV sector A0 service building. Operations contacted the cryo system technicians at CHL. The TeV sector F0 cold compressor was brought online.

Operations established store 5262 at 10:06 AM with an initial luminosity of 216.42E30.

At 10:13 AM, while operators were trying to turn off the Pelletron after the shot, they noticed “DPM_Pend” reports on all Electron Cooling devices. Operations contacted experts, accessed the Pelletron enclosure, rebooted its VME locally, and had everything operating normally within two hours.

At 3:53 PM, Operations reported a problem with the Switchyard (SY) vacuum degrading. Operations contacted experts.

Operations reported at 7:59 PM that the Recycler had stashed four times during the last 12 hours.

At 11:57 PM, MI tripped off due to the MI-8 collimators. It reset.

Monday March 5

The midnight shift began with Operations monitoring store 5262 and the stash, with Pbar stacking, and with NuMI taking beam.

At 1:28 AM, NuMI tripped off due to QD108 current drift. It reset.

At 3:16 AM, the Pbar target stopped rotating. Operations contacted an expert. Per instruction, operators reversed the direction of the target at 4:04 AM.

At 5 AM, the Pbar target stopped rotating again. Operators turned off Pbar beam in preparation for an access later in the morning.

SY experts began taking study beam at 7 AM.

Operations terminated store 5262 at 8:01 AM.

Operations reported at 8:10 AM that the Recycler had stashed four times during the last 12 hours.

The Plans for Monday

The plans for today are to allow CDF and D0 experimenters to access their collision halls. Pbar experts will repair their target.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/05/2007    email Fermilab